OpenStudy (anonymous):

A company is considering making a new product. They estimate the probability that the new product will be successful is 0.75. If it is successful it would generate $240,000 in profit. The cost to develop the product is $196,000. Use the revenue (profit – cost) and expected value to decide whether the company should make this new product.

OpenStudy (amistre64):

what are your thoughts for a solution?

OpenStudy (anonymous):

I am not really sure how to do it~

OpenStudy (amistre64):

you take the probability of success (and the revenue generated by it) and the probability of failure (and the cost associated with trying it) and get an expected value aP(a) + bP(b)

OpenStudy (anonymous):

Okay

OpenStudy (amistre64):

spose you have a 1/4 chance of winning 3 dollars, and all you have to do is pay 1 dollar your expected winning is just: 3(1/4) - 1(3/4) = 0, you expect to gain 0 dollars.

OpenStudy (amistre64):

what do you get from your problem?

OpenStudy (anonymous):

0.27(150,000) + 0.73(–35,000) = E(X) ?

OpenStudy (anonymous):

or is it this 0.27(115,000) + 0.73(–35,000) = E(X)? I think this more makes

OpenStudy (amistre64):

cost is a negative since we lose money; success is .75 so failure is .25 .75(revenue) - .25(cost) .75(240000-196000) - .25(196000) .75(44) - .25(196)

OpenStudy (anonymous):

0.27(150,000) + 0.73(–35,000) = E(X) 150,000 – 0.73(35,000) = E(X) 0.27(150,000 – 35,000) = E(X) 0.27(115,000) + 0.73(–35,000) = E(X) So would it be the last one?

OpenStudy (anonymous):

anyone there I really need help

OpenStudy (anonymous):

Also I have this one too Two students created a game of dice. They determined that each roll would add the value of the die if it was an odd number. If the roll was even, it would subtract the value of the die. (For example, a roll of 2 would subtract two points, but a roll of 3 would add three points.) What is the expected value for this game? –3 –0.5 0 0.5 I believe the answer is -0.5 but don't know

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

@brian16

OpenStudy (amistre64):

your options on your first do not seem to correspond to the information given

OpenStudy (amistre64):

each die face has a 1/6 probability assuming its a 6faced die -1(1/6) -3(1/6) -5(1/6) +2(1/6) +4(1/6) +6(1/6) would give us the expected score of the game if the die is rolled once

OpenStudy (amistre64):

well, i got my odds and evens mixed up ....

OpenStudy (amistre64):

and once again, your options do not correspond to the information ....

OpenStudy (anonymous):

okay so would it be 0.27(150,000) + 0.73(–35,000) = E(X)

OpenStudy (amistre64):

lets see if i remember how to add 10 and 12 ... 2/6 = 1/3

OpenStudy (amistre64):

.75 and .25 are not .73 and .27 your information does not correspond to your options. either fix the question, or tell a teacher that the question is broken.

OpenStudy (anonymous):

Okay thanks

OpenStudy (anonymous):

these are the other things i would think 150,000 – 0.73(35,000) = E(X) 0.27(150,000 – 35,000) = E(X)

OpenStudy (anonymous):

which one do you think?

OpenStudy (anonymous):

I really need HELP!!!

OpenStudy (anonymous):

AS u read carefully it says it can make 240 000 profit and 196 000 loss otherwise so i would think: 0.75*240-0.25*196

OpenStudy (anonymous):

thank you so much!!!!!

OpenStudy (anonymous):

I GOT IT RIGHT YOUR AMAZING!!!